OpenStudy (anonymous):

Hi, I am wondering whether the sequence cos nπ/2 converges, diverges, or alternates between -1 and 1? thanks

OpenStudy (anonymous):

It alternates.

OpenStudy (anonymous):

how about this one? thanks [-1+((-1)^{n ^{2}}/ n ^{2})\]

OpenStudy (anonymous):

what happens as n gets really, really large?

OpenStudy (anonymous):

an gets close negative 1

OpenStudy (anonymous):

yup. there you go.

OpenStudy (anonymous):

so diverges eh?

OpenStudy (anonymous):

uh, no... it converges to -1...

OpenStudy (anonymous):

so if n gets larger, and an approaches 0 or a number, the series converge?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks a lot mate

OpenStudy (anonymous):

sorry to disturb again, I thought that if lim n-infinity, and an approaches only 0, it is a convergence series, and if approaches infinity or any real number it is a dvergence series, Is there any exception regarding this theory?

OpenStudy (anonymous):

Oh, I'm sorry. I misunderstood your question. The sequence converges if \[ \lim_{n\rightarrow \infty} a_n = L\] where L is any real number. The convergence of a SERIES is more complex. If a series is to converge, then \[ \lim_{n\rightarrow \infty} a_n = 0 \] but this is not the only requirement.

OpenStudy (anonymous):

okay, last thing what is the requirement for a sequence ( not of a series) to diverge?? thank a lot

OpenStudy (anonymous):

A sequence diverges if the limit as n -> infinity doesn't exist. Like, \[ a_n = \log(n) \]